# Approximations for the number of maxima and near-maxima in independent data

### v1

@@ 2 Stein’s method for the logarithmic distribution
$\mathbb{E}f(L^{\star}-1)$ $=(1-\alpha)\sum_{k=0}^{\infty}\alpha^{k}f(k)=(1-\alpha)f(0)-(1-\alpha)\log(1-\alpha)\sum_{k=1}^{\infty}kf(k)\mathbb{P}(L=k)$
@@
=$\mathcal{H}=\{I(\cdot\in E)-\mathbb{P}(L\in E):E\subseteq\mathbb{Z}^{+}\}$ (9)
@@
$f_{h}(k)$ $=-\frac{1}{\alpha}\sum_{j=1}^{k}h(j)\frac{\alpha^{j-1}}{j}=\frac{\log(1-\alpha)}{\alpha^{k+1}}\sum_{j=1}^{k}h(j)\mathbb{P}(L=j)$
$=-\frac{\log(1-\alpha)}{\alpha^{k+1}}\sum_{j=k+1}^{\infty}h(j)\mathbb{P}(L=j)=\frac{1}{\alpha}\sum_{j=k+1}^{\infty}h(j)\frac{\alpha^{j-k}}{j}=\frac{1}{\alpha}\sum_{j=1}^{\infty}h(j+k)\frac{\alpha^{j}}{j+k}\,,$
@@ 2.2 Proof of Theorem 1(a)
To establish Theorem 1(a), we combine the mass function and moments given in (2) and (3), respectively, with the following result.
@@
$\mathbb{P}(\widetilde{K}^{\star}-1=K^{\star}|\widetilde{K}^{\star}>1)$ $\geq\mathbb{P}(\widetilde{K}^{\star}=2,K^{\star}=1|\widetilde{K}^{\star}>1)=\mathbb{P}(\widetilde{K}^{\star}=2|\widetilde{K}^{\star}>1)\mathbb{P}(K^{\star}=1)$
@@
When bounding the probability $\mathbb{P}(\widetilde{K}^{\star}-1=K^{\star}|\widetilde{K}^{\star}>1)$ in (2.2) we could, of course, include further terms of the form $\mathbb{P}(\widetilde{K}^{\star}=j+1,K^{\star}=j|\widetilde{K}^{\star}>1)$ for $j>1$ to get a more precise lower bound. However, in our coupling construction we have in mind the setting in which $\alpha$ is small, so that both $\widetilde{K}^{\star}-1$ and $I_{\alpha}K^{\star}$ are equal to zero with high probability. In this setting, further terms of this sort for $j>1$ are significantly smaller than the $j=1$ term we have included in the bound (2.2), so add complexity to the resulting bound on total variation distance for no significant gain in accuracy.
@@ 3 Negative binomial approximation for mixed binomials
Our main aim in this section is to prove Theorems 1(b) and 5. To do this, we begin by considering the more general setting of negative binomial approximation for mixed binomial random variables. We let $W$ have a mixed binomial distribution. In particular, for some $n\in\{1,2,\ldots\}$, $\ell\in\{1,\ldots,n\}$ and for some random variable $Q$ supported on $[0,1]$, we let the conditional distribution $W|Q\sim\text{Bin}(n-\ell,Q)$ have a binomial distribution with parameters $n-\ell$ and $Q$. We then write that $W\sim\text{MBin}(n-\ell,Q)$. We consider the approximation of $W$ by the negative binomial random variable $Z\sim\text{NB}(\ell,1-\beta)$, where we choose $\beta=\frac{\mathbb{E}[W]}{\mathbb{E}[W]+\ell}$ so that $\mathbb{E}[Z]=\mathbb{E}[W]$. Our main result here is the following.
@@
$=\frac{(1-\beta)(\ell-1)}{\ell}\mathbb{E}[g_{h}(W+1)]+\left(1-\frac{(1-\beta)(\ell-1)}{\ell}\right)\mathbb{E}[g_{h}((W+1)^{\star})]\,,$
where the final equality uses the definition (6) of size-biasing. Again using this same definition, we may therefore write
@@
where $W^{\prime}$ is equal in distribution to $I_{\gamma}(W+1)^{\star}+(1-I_{\gamma})(W+1)$ and $I_{\gamma}$ is a Bernoulli random variable with mean $\gamma=1-\frac{(1-\beta)(\ell-1)}{\ell}$ independent of all other random variables with which it appears.
It is well-known that we may size-bias a sum of independent random variables by selecting one of these random variables with probabilities proportional to their means, and replacing the chosen random variable with its size-biased version; see Section 2.4 of [1]. Combining this with Lemma 2.4 of [1] on size-biasing a mixture, we have that $W^{\star}=1+A$, where $A\sim\text{MBin}(n-\ell-1,Q^{\star})$. Similarly, and noting that $\frac{\mathbb{E}[W]}{\mathbb{E}[W]+1}=\frac{\beta}{\gamma}$, we may write
@@
$W^{\prime}=1+\left[1-I_{\gamma}I_{\beta/\gamma}\right]W+I_{\gamma}I_{\beta/\gamma}W^{\star}=1+I_{\gamma}I_{\beta/\gamma}+B+C\,,$
@@
$=\frac{1-(1-\beta)^{\ell}}{\beta\ell}\left(\beta+(1-\beta)\left\{(n-\ell-1)\mathbb{E}[Q^{\star}]-(n-\ell-2)\mathbb{E}[Q]\right\}\right)\,.$ (17)
@@ 3.1 Proof of Theorem 1(b)
From the definition of the geometric distribution we can easily check that $G-1$ is equal in distribution to $I_{\beta}G$. We may thus use the triangle inequality for total variation distance to write
$d_{\text{TV}}(K^{\star}-1,I_{\beta}K^{\star})$ $\leq d_{\text{TV}}(K^{\star}-1,G-1)+d_{\text{TV}}(I_{\beta}K^{\star},I_{\beta}G)=d_{\text{TV}}(K^{\star}-1,G-1)+\beta d_{\text{TV}}(K^{\star},G)$
@@
  • Set $K_{n}^{\star}-1=|\left\{i\in\{1,\ldots,n-1\}:X_{i}^{\prime}=M\right\}|$
@@
$d_{\text{TV}}(K,L)\\ \leq-\frac{2(1+\beta)\log(1-\beta)}{\beta}\mathbb{E}[K_{n}]\mathbb{E}[K_{n}^{\star}-1]\left(\beta+(1-\beta)\left[(n-2)\frac{\mathbb{E}[q(M)^{2}]}{\mathbb{E}[q(M)]}-(n-3)\mathbb{E}[q(M)]\right]\right)\,.$
@@ 4 Poisson approximation for $K_{n}$ in the discrete case
$d_{\text{TV}}(K_{n}^{\star}-1,Y)\leq\mathbb{E}[q(M)]+\frac{(n-1)^{2}\text{Var}(q(M))}{\lambda}=\frac{(n-1)\mathbb{E}[(K_{n})_{3}]}{(n-2)\mathbb{E}[(K_{n})_{2}]}-\frac{(n-2)\mathbb{E}[(K_{n})_{2}]}{(n-1)\mathbb{E}[K_{n}]}\,,$ (24)

### v2

@@ 2 Stein’s method for the logarithmic distribution
$\mathbb{E}f(L^{\star}-1)$ $=(1-\alpha)\sum_{k=0}^{\infty}\alpha^{k}f(k)$
$=(1-\alpha)f(0)-(1-\alpha)\log(1-\alpha)\sum_{k=1}^{\infty}kf(k)\mathbb{P}(L=k)$
@@
$\mathcal{H}=\{I(\cdot\in E)-\mathbb{P}(L\in E):E\subseteq\mathbb{Z}^{+}\}$ (9)
@@
$f_{h}(k)$ $=-\frac{1}{\alpha^{k+1}}\sum_{j=1}^{k}h(j)\frac{\alpha^{j}}{j}=\frac{\log(1-\alpha)}{\alpha^{k+1}}\sum_{j=1}^{k}h(j)\mathbb{P}(L=j)$
$=-\frac{\log(1-\alpha)}{\alpha^{k+1}}\sum_{j=k+1}^{\infty}h(j)\mathbb{P}(L=j)=\frac{1}{\alpha}\sum_{j=k+1}^{\infty}h(j)\frac{\alpha^{j-k}}{j}$
$=\frac{1}{\alpha}\sum_{j=1}^{\infty}h(j+k)\frac{\alpha^{j}}{j+k}\,$,
@@ 2.2 Proof of Theorem 1(a)
To establish Theorem 1(a), we combine the mass function and moments given in (2) and (3), respectively, with the following result giving an upper bound on the total variation distance between a positive, integer-valued random variable $K$ and $L\sim\text{L}(\alpha)$, where we choose $\alpha=\mathbb{P}(K^{\star}>1)$. This choice is convenient in the proof (as we will see following (11) below), and also gives
$1-\alpha=\frac{\mathbb{P}(K=1)}{\mathbb{E}[K]}=\frac{\mathbb{P}(L=1)}{\mathbb{E}[L]}\,.$
@@
$\mathbb{P}(\widetilde{K}^{\star}-1=K^{\star}|\widetilde{K}^{\star}>1)$ $\geq\mathbb{P}(\widetilde{K}^{\star}=2,K^{\star}=1|\widetilde{K}^{\star}>1)$
$=\mathbb{P}(\widetilde{K}^{\star}=2|\widetilde{K}^{\star}>1)\mathbb{P}(K^{\star}=1)$
@@
When bounding the probability $\mathbb{P}(\widetilde{K}^{\star}-1=K^{\star}|\widetilde{K}^{\star}>1)$ in (12) we could, of course, include further terms of the form $\mathbb{P}(\widetilde{K}^{\star}=j+1,K^{\star}=j|\widetilde{K}^{\star}>1)$ for $j>1$ to get a more precise lower bound. However, in our coupling construction we have in mind the setting in which $\alpha$ is small, so that both $\widetilde{K}^{\star}-1$ and $I_{\alpha}K^{\star}$ are equal to zero with high probability. In this setting, further terms of this sort for $j>1$ are significantly smaller than the $j=1$ term we have included in the bound (12), so add complexity to the resulting bound on total variation distance for no significant gain in accuracy. Related to this, we note that that even if $K$ is logarithmically distributed, the upper bound of Theorem 9 is not zero. Using $\mathbb{E}[L]=\frac{-\alpha}{(1-\alpha)\log(1-\alpha)}$ and $\mathbb{P}(L=2)=\frac{-\alpha^{2}}{2\log(1-\alpha)}$, the upper bound of Theorem 9 with $K=L$ is equal to
$2\alpha\left(\frac{1}{1-\alpha}-(1-\alpha)\right)\,$, (14)
which is increasing in $\alpha$.
@@ 3 Negative binomial approximation for mixed binomials
Our main aim in this section is to prove Theorems 1(b) and 5. To do this, we begin by considering the more general setting of negative binomial approximation for mixed binomial random variables. We let $W$ have a mixed binomial distribution. In particular, for some $n\in\{1,2,\ldots\}$, $\ell\in\{1,\ldots,n-1\}$ and for some random variable $Q$ supported on $[0,1]$, we let the conditional distribution $W|Q\sim\text{Bin}(n-\ell,Q)$ have a binomial distribution with parameters $n$-$\ell$ and $Q$. We then write that $W\sim\text{MBin}(n-\ell,Q)$. We consider the approximation of $W$ by the negative binomial random variable $Z\sim\text{NB}(\ell,1-\beta)$, where we choose $\beta=\frac{\mathbb{E}[W]}{\mathbb{E}[W]+\ell}$ so that $\mathbb{E}[Z]=\mathbb{E}[W]$. Our main result here is the following.
@@
$=\frac{(1-\beta)(\ell-1)}{\ell}\mathbb{E}[g_{h}(W+1)]$
$\qquad\qquad\qquad+\left(1-\frac{(1-\beta)(\ell-1)}{\ell}\right)\mathbb{E}[g_{h}((W+1)^{\star})]\,$,
where the final equality uses the definition (6) of size-biasing. Hence,
$\mathbb{E}[\beta(\ell+W)g_{h}(W+1)]$ $=\mathbb{E}[W]\left\{(1-\gamma)\mathbb{E}[g_{h}(W+1)]+\gamma\mathbb{E}[g_{h}((W+1)^{\star})]\right\}$
$=\mathbb{E}[W]\mathbb{E}\left[g_{h}((1-I_{\gamma})(W+1)+I_{\gamma}(W+1)^{\star})\right]\,$,
where $\gamma=1-\frac{(1-\beta)(\ell-1)}{\ell}$ and $I_{\gamma}$ is a Bernoulli random variable with mean $\gamma$ independent of all other random variables with which it appears. Again using the definition of size-biasing, we may therefore write
@@
where $W^{\prime}$ is equal in distribution to $(1-I_{\gamma})(W+1)+I_{\gamma}(W+1)^{\star}$.
It is well-known that we may size-bias a sum of independent random variables by selecting one of these random variables with probabilities proportional to their means, and replacing the chosen random variable with its size-biased version; see Section 2.4 of [1]. Combining this with Lemma 2.4 of [1] on size-biasing a mixture, we have that $W^{\star}=1+A$, where $A\sim\text{MBin}(n-\ell-1,Q^{\star})$ and a mixed binomial distribution with first parameter equal to zero should be interpreted as a point mass at zero. Similarly, and noting that $\frac{\mathbb{E}[W]}{\mathbb{E}[W]+1}=\frac{\beta}{\gamma}$, we may write
@@
$W^{\prime}=1+\left[1-I_{\gamma}I_{\beta/\gamma}\right]W+I_{\gamma}I_{\beta/\gamma}W^{\star}=1+I_{\gamma}I_{\beta/\gamma}+B+C\,$,\
@@
$=\frac{1-(1-\beta)^{\ell}}{\beta\ell}(\beta+(1-\beta)\{(n-\ell-1)\mathbb{E}[Q^{\star}]-(n-\ell-2)\mathbb{E}[Q]\})\,.$ (18)
@@ 3.1 Proof of Theorem 1(b)
From the definition of the geometric distribution we can easily check that G1G-1 is equal in distribution to $I_{\beta}G$. We may thus use the triangle inequality for total variation distance to write
$d_{\text{TV}}(K^{\star}-1,I_{\beta}K^{\star})$ $\leq d_{\text{TV}}(K^{\star}-1,G-1)+d_{\text{TV}}(I_{\beta}K^{\star},I_{\beta}G)$
$=d_{\text{TV}}(K^{\star}-1,G-1)+\beta d_{\text{TV}}(K^{\star},G)$
@@
  • Set $K_{n}^{\star}-1=|\left\{i\in\{1,\ldots,n-1\}:X_{i}^{\prime}=M\right\}|$ .
@@
$d_{\text{TV}}(K,L)\leq-\frac{2(1+\beta)\log(1-\beta)}{\beta}\mathbb{E}[K_{n}]\mathbb{E}[K_{n}^{\star}-1]\\ \times\left(\beta+(1-\beta)\left[(n-2)\frac{\mathbb{E}[q(M)^{2}]}{\mathbb{E}[q(M)]}-(n-3)\mathbb{E}[q(M)]\right]\right)\,.$
@@
Remark 13.
By considering how the upper bound of Theorem 1(b) behaves should the first three moments of $K_{n}$ match those of a logarithmic distribution, we would expect this to be generally inferior to Theorem 1(a). Using the fact that
$\mathbb{E}[(L)_{\ell}]=\frac{-(\ell-1)!}{\log(1-\alpha)}\left(\frac{\alpha}{1-\alpha}\right)^{\ell}$
for $L\sim\text{L}(\alpha)$ and $\ell=1,2,\ldots$, with $\beta=\alpha$ and the factorial moments of $K_{n}$ replaced by those of $L$, the upper bound of Theorem 1(b) becomes
$\frac{2\alpha(1+\alpha)}{(1-\alpha)^{2}}\left(\alpha+(1-\alpha)\left[\frac{2\alpha}{1-\alpha}-\frac{(n-3)\alpha}{(n-1)(1-\alpha)}\right]\right)\,$,
which for large $n$ is approximately equal to $\frac{4\alpha^{2}(1+\alpha)}{(1-\alpha)^{2}}$. For all $\alpha\in(0,1)$ this is greater than (14), which gives a comparable expression for the bound of Theorem 1(a). Although we typically expect that part (a) of Theorem 1 is superior to part (b), we nevertheless include both since we have been unable to show that this is the case for all possible choices of the underlying distribution of the $X_{i}$.
@@ 4 Poisson approximation for $K_{n}$ in the discrete case
$d_{\text{TV}}(K_{n}^{\star}-1,Y)$ $\leq\mathbb{E}[q(M)]+\frac{(n-1)^{2}\text{Var}(q(M))}{\lambda}$
$=\frac{(n-1)\mathbb{E}[(K_{n})_{3}]}{(n-2)\mathbb{E}[(K_{n})_{2}]}-\frac{(n-2)\mathbb{E}[(K_{n})_{2}]}{(n-1)\mathbb{E}[K_{n}]}\,$, (25)
@@
5 Concluding remarks
In this note we have established explicit error bounds in the logarithmic and Poisson approximation of the number of maxima in a sample of independent, discrete data. The setting in which the underlying data are geometrically distributed was our main example here. Similarly, in the case of continuous data, illustrated by the Gumbel and uniform examples, we establish error bounds in negative binomial approximation for the number of data points within a given threshold of one of the order statistics of our sample.
We conclude by noting scope for further work. We expect that, with a more sophisticated coupling than used here, error bounds and rates of convergence could be strengthened; for example we would expect to be able to obtain an error bound which decays to zero for a fixed $a$ in the setting of Example 6. This may require allowing both parameters of our approximating negative binomial distribution $Z$ in Theorem 11 to depend on $Q$, rather than only the second parameter. In other cases we similarly note that there is scope to find informative error bounds for a greater range of underlying parameter values. For example, in the geometric setting of Example 2 it remains an open problem to find reasonable error bounds for larger values of the parameter $p$. Our geometric and Gumbel examples also illustrate the scope to improve the tightness of our upper bounds.
It would be of interest to extend our findings beyond the case of independent data. We may then no longer have explicit expressions for the distribution of $K_{n}$ or the order statistics of our data, and so we would expect proofs to become more complex, but note that an important general advantage Stein’s method over other techniques to derive such error bounds is its ability to cope with relaxation of independence assumptions.
Finally, we note that the development of Stein’s method for approximation by a logarithmic distribution was an important step in our work. In future work we hope to be able to make use of these tools in applications beyond the setting of the present paper.
Acknowledgements
The author thanks two anonymous reviewers and an Editor for their careful readings of an earlier version of this manuscript, and constructive comments and suggestions that improved the paper.